Chapter 3

Discrete random variables

[144/99%]
Show
LaTeX
Problem 3.1.1

For what values of the constant CC do the following define mass functions on the positive integers 1,2,…1,2, \ldots?

?
(a)

Geometric: f(x)=C2−xf(x) = C 2^{-x}.

(b)

Logarithmic: f(x)=C2−x/xf(x) = C 2^{-x} / x.

(c)

Inverse square: f(x)=Cx−2f(x) = C x^{-2}.

(d)

'Modified' Poisson: f(x)=C2x/xf(x) = C 2^{x} / x !.

Problem 3.1.2

For a random variable XX having each of the four mass functions of Exercise (3.1.1), find P(X>1)\mathbb {P}\left(X > 1\right), the most probable value of XX, and the probability that XX is even, when the mass function is:

?
(a)

Geometric (Exercise 3.1.1(a)): f(x)=C2−xf(x) = C 2^{-x}.

(b)

Logarithmic (Exercise 3.1.1(b)): f(x)=C2−x/xf(x) = C 2^{-x} / x.

(c)

Inverse square (Exercise 3.1.1(c)): f(x)=Cx−2f(x) = C x^{-2}.

(d)

'Modified' Poisson (Exercise 3.1.1(d)): f(x)=C2x/x!f(x) = C 2^{x} / x!.

Problem 3.1.3

We toss nn coins, and each one shows heads with probability pp, independently of each of the others. Each coin which shows heads is tossed again. What is the mass function of the number of heads resulting from the second round of tosses?

?
Problem 3.1.4

Let SkS_{k} be the set of positive integers whose base-10 expansion contains exactly kk elements (so that, for example, 1024∈S41024 \in S_{4}). A fair coin is tossed until the first head appears, and we write TT for the number of tosses required. We pick a random element, NN say, from STS_{T}, each such element having equal probability. What is the mass function of NN?

?
Problem 3.1.5
?
(a)

Show that, if XX is a binomial or Poisson random variable, then the mass function f(k)=P(X=k)f(k) = \mathbb {P}\left(X = k\right) has the property that f(k−1)f(k+1)≤f(k)2f(k-1) f(k+1) \leq f(k)^{2}.

(b)

Show that, if f(k)=90/(πk)4,k≥1f(k) = 90 /(\pi k)^{4}, k \geq 1, then f(k−1)f(k+1)≥f(k)2f(k-1) f(k+1) \geq f(k)^{2}.

(c)

Find a mass function ff such that f(k)2=f(k−1)f(k+1),k≥1f(k)^{2} = f(k-1) f(k+1), k \geq 1.

Problem 3.2.1

Let XX and YY be independent random variables, each taking the values −1-1 or 11 with probability 12\frac{1}{2}, and let Z=XYZ = X Y. Show that X,YX, Y, and ZZ are pairwise independent. Are they independent?

?
Problem 3.2.2

Let XX and YY be independent random variables taking values in the positive integers and having the same mass function f(x)=2−xf(x) = 2^{-x} for x=1,2,…x = 1,2, \ldots. Find:

?
(a)

P(min⁡{X,Y}≤x)\mathbb {P}\left(\min \left\{ X, Y\right\} \leq x\right),

(b)

P(Y>X)\mathbb {P}\left(Y > X\right),

(c)

P(X=Y)\mathbb {P}\left(X = Y\right),

(d)

P(X≥kY)\mathbb {P}\left(X \geq k Y\right), for a given positive integer kk,

(e)

P(X divides Y)\mathbb {P}\left(X\text{ divides }Y\right),

(f)

P(X=rY)\mathbb {P}\left(X = r Y\right), for a given positive rational rr.

Problem 3.2.3

Let X1,X2,X3X_{1}, X_{2}, X_{3} be independent random variables taking values in the positive integers and having mass functions given by P(Xi=x)=(1−pi)pix−1\mathbb {P}\left(X_{i} = x\right) = \left(1-p_{i}\right) p_{i}^{x-1} for x=1,2,…x = 1,2, \ldots, and i=1,2,3i = 1,2,3.

?
(a)

Show that

P(X1<X2<X3)=(1−p1)(1−p2)p2p32(1−p2p3)(1−p1p2p3) \mathbb {P}\left(X_{1} < X_{2} < X_{3}\right) = \frac{\left(1-p_{1}\right)\left(1-p_{2}\right) p_{2} p_{3}^{2}}{\left(1-p_{2} p_{3}\right)\left(1-p_{1} p_{2} p_{3}\right)}
(b)

Find P(X1≤X2≤X3)\mathbb {P}\left(X_{1} \leq X_{2} \leq X_{3}\right).

Problem 3.2.4

Three players, A,B\mathrm{A}, \mathrm{B}, and C, take turns to roll a die; they do this in the order ABCABCA....

?
(a)

Show that the probability that, of the three players, AA is the first to throw a 66, BB the second, and C the third, is 216/1001216 / 1001.

(b)

Show that the probability that the first 6 to appear is thrown by A, the second 6 to appear is thrown by B, and the third 6 to appear is thrown by C, is 46656/75357146656 / 753571.

Problem 3.2.5

Let Xr,1≤r≤nX_{r}, 1 \leq r \leq n, be independent random variables which are symmetric about 0 ; that is, XrX_{r} and −Xr-X_{r} have the same distributions. Show that, for all x,P(Sn≥x)=P(Sn≤−x)x, \mathbb {P}\left(S_{n} \geq x\right) = \mathbb {P}\left(S_{n} \leq -x\right) where Sn=∑r=1nXrS_{n} = \sum_{r = 1}^{n} X_{r}.

Is the conclusion necessarily true without the assumption of independence?

?
Problem 3.3.1
?
(a)

Is it generally true that E[1/X]=1/E[X]\mathbb {E}\left[1 / X\right] = 1 / \mathbb {E}\left[X\right]?

(b)

Is it ever true that E[1/X]=1/E[X]\mathbb {E}\left[1 / X\right] = 1 / \mathbb {E}\left[X\right]?

Problem 3.3.2

Every package of some intrinsically dull commodity includes a small and exciting plastic object. There are cc different types of object, and each package is equally likely to contain any given type. You buy one package each day.

?
(a)

Find the mean number of days which elapse between the acquisitions of the jj th new type of object and the (j+1)(j+1) th new type.

(b)

Find the mean number of days which elapse before you have a full set of objects.

Problem 3.3.3

Each member of a group of nn players rolls a die.

?
(a)

For any pair of players who throw the same number, the group scores 1 point. Find the mean and variance of the total score of the group.

(b)

Find the mean and variance of the total score if any pair of players who throw the same number scores that number.

Problem 3.3.4

A fair coin is tossed repeatedly. Let TT be the number of tosses until the first head. You are offered the following prospect, which you may accept on payment of a fee. If T=kT = k, say, then you will receive £2k£ 2^{k}. What would be a 'fair' fee to ask of you?

?
Remark.
?

The use of expectation to assess a 'fair fee' may be convenient but is sometimes inappropriate. For example, a more suitable criterion in the finance market would be absence of arbitrage; see Exercise (3.3.7) and Section 13.10. And, in a rather general model of financial markets, there is a criterion commonly expressed as 'no free lunch with vanishing risk'.

Problem 3.3.5

Let XX have mass function

f(x)={(x(x+1))−1 if x=1,2,…0 otherwise  f(x) = \begin{cases} \left(x(x+1)\right)^{-1} & \text{ if } x = 1,2, \ldots \\ 0 & \text{ otherwise }\end{cases}

and let α∈R\alpha \in \mathbb {R}. For what values of α\alpha is it the case that E[Xα]<∞\mathbb {E}\left[X^{\alpha }\right] < \infty?

?
Problem 3.3.6

Show that Var⁡[a+X]=Var⁡[X]\operatorname {Var}\left[a+X\right] = \operatorname {Var}\left[X\right] for any random variable XX and constant aa.

?
Problem 3.3.7

Suppose you find a warm-hearted bookmaker offering payoff odds of π(k)\pi (k) against the kkth horse in an nn-horse race where ∑k=1n(π(k)+1)−1<1\sum_{k = 1}^{n}\left(\pi (k)+1\right)^{-1} < 1. Show that you can distribute your bets in such a way as to ensure you win.

?
Remark.
?

If one bets on the kkth horse xx money, he gets (1+π(k))x(1 + \pi (k))x if kkth horse wins nothing otherwise.

Problem 3.3.8

You roll a conventional fair die repeatedly. If it shows 11, you must stop, but you may choose to stop at any prior time. Your score is the number shown by the die on the final roll.

?
(a)

What stopping strategy yields the greatest expected score?

(b)

What strategy would you use if your score were the square of the final roll?

Problem 3.3.9

Continuing with Exercise (3.3.8), suppose now that you lose cc points from your score each time you roll the die. What strategy maximizes the expected final score if c=13c = \frac{1}{3}? What is the best strategy if c=1c = 1?

?
Problem 3.3.10

Let G=(V,E)G = (V, E) be a random graph with m=∣V∣m = \left|V\right| vertices and edge-set EE. Write dvd_{v} for the degree of vertex vv, that is, the number of edges meeting at vv. Let YY be a uniformly chosen vertex, and ZZ a uniformly chosen neighbour of YY.

?
(a)

Show that E[dZ]≥E[dY]\mathbb {E}\left[d_{Z}\right] \geq \mathbb {E}\left[d_{Y}\right].

(b)

Interpret this inequality when the vertices represent people, and the edges represent friendship.

Problem 3.3.11

The gambler Lester Savage makes up to three successive bets that a fair coin flip will show heads. He places a stake on each bet, which, if a head is shown, pays him back twice the stake. If a tail is shown, he loses his stake.

His stakes are determined as follows. Let x>y>z>0x > y > z > 0. He bets xx on the first flip; if it shows heads he quits, otherwise he continues. If he continues, he bets yy on the second flip; if it shows heads he quits, otherwise he continues. If he continues, he bets zz on the third flip.

Let GG be his accumulated gain (positive or negative). List the possible values of GG and their probabilities. Show that E[G]=0\mathbb {E}\left[G\right] = 0, and find Var⁡[G]\operatorname {Var}\left[G\right] and P(G<0)\mathbb {P}\left(G < 0\right).

Lester decides to stick with the three numbers x,y,zx, y, z but to vary their order. How should he place his bets in order to simultaneously minimize both P(G<0)\mathbb {P}\left(G < 0\right) and Var⁡[G]\operatorname {Var}\left[G\right]? Explain.

?
Problem 3.3.12

A set of nn different words is equally likely to be in any of the nn ! possible orders. It is decided to place them in lexicographic order using the following algorithm.

(i) Compare the first word ww with the others, and find the set of earlier words and the set of later words.

(ii) Iterate the procedure on each of the two sets thus obtained.

(iii) Continue until the final ordering is achieved.

?
(a)

Give an expression for the mean number cnc_{n} of comparisons required.

(b)

Show that cn=2n(log⁡n+γ−2)+O(1)c_{n} = 2 n(\log n+\gamma -2)+\mathrm{O}(1) as n→∞n \rightarrow \infty, where γ\gamma is Euler's constant.

(c)

Let nn be replaced by a random variable NN with mass function

P(N=n)=A(n−1)n(n+1),n≥2 \mathbb {P}\left(N = n\right) = \frac{A}{(n-1) n(n+1)}, \quad n \geq 2

for suitable AA. Show that the mean number of comparisons is 4.

Problem 3.4.1
?
(a)

A biased coin is tossed nn times, and heads shows with probability pp on each toss. A run is a sequence of throws which result in the same outcome, so that, for example, the sequence HHTHTTH contains five runs. Show that the expected number of runs is 1+2(n−1)p(1−p)1+2(n-1) p(1-p). Find the variance of the number of runs.

(b)

Let hh heads and tt tails be arranged randomly in a line. Find the mean and variance of the number of runs of heads

Problem 3.4.2

An urn contains nn balls numbered 1,2,…,n1,2, \ldots , n. We remove kk balls at random (without replacement) and add up their numbers. Find the mean and variance of the total.

?
Problem 3.4.3

Of the 2n2 n people in a given collection of nn couples, exactly mm die. Assuming that the mm have been picked at random, find the mean number of surviving couples. This problem was formulated by Daniel Bernoulli in 1768.

?
Problem 3.4.4

Urn R contains nn red balls and urn B contains nn blue balls. At each stage, a ball is selected at random from each urn, and they are swapped. Show that the mean number of red balls in urn RR after stage kk is 12n{1+(1−2/n)k}\frac{1}{2} n\left\{ 1+(1-2 / n)^{k}\right\}. This 'diffusion model' was described by Daniel Bernoulli in 1769.

?
Problem 3.4.5

Consider a square with diagonals, with distinct source and sink. Each edge represents a component which is working correctly with probability pp, independently of all other components. Write down an expression for the Boolean function which equals 1 if and only if there is a working path from source to sink, in terms of the indicator functions XiX_{i} of the events {edge i is working}\left\{ \text{edge }i\text{ is working}\right\} as ii runs over the set of edges. Hence calculate the reliability of the network.

?
Problem 3.4.6

A system is called a ' kk out of nn ' system if it contains nn components and it works whenever kk or more of these components are working. Suppose that each component is working with probability pp, independently of the other components, and let XcX_{c} be the indicator function of the event that component cc is working. Find, in terms of the XcX_{c}, the indicator function of the event that the system works, and deduce the reliability of the system.

?
Problem 3.4.7

Let G=(V,E)G = (V, E) be a finite graph. For any set WW of vertices and any edge e∈Ee \in E, define the indicator function

IW(e)={1 if e connects W and W∁0 otherwise  I_{W}(e) = \begin{cases} 1 & \text{ if } e \text{ connects } W \text{ and } W^{\complement } \\ 0 & \text{ otherwise }\end{cases}

Set NW=∑e∈EIW(e)N_{W} = \sum_{e \in E} I_{W}(e). Show that there exists W⊆VW \subseteq V such that NW≥12∣E∣N_{W} \geq \frac{1}{2}\left|E\right|.

?
Problem 3.4.8

A total of nn bar magnets are placed end to end in a line with random independent orientations. Adjacent like poles repel, ends with opposite polarities join to form blocks. Let XX be the number of blocks of joined magnets. Find E[X]\mathbb {E}\left[X\right] and Var⁡[X]\operatorname {Var}\left[X\right].

?
Problem 3.4.9
?
(a)

Use the inclusion-exclusion formula (3.4.2) to derive the result of Example (3.4.3), namely: in a random permutation of the first nn integers, the probability that exactly rr retain their original positions is

1r!(12!−13!+⋯+(−1)n−r(n−r)!) \frac{1}{r!}\left(\frac{1}{2!}-\frac{1}{3!}+\cdots +\frac{(-1)^{n-r}}{(n-r)!}\right)
(b)

Let dnd_{n} be the number of derangements of the first nn integers (that is, rearrangements with no integers in their original positions). Show that dn+1=ndn+ndn−1d_{n+1} = n d_{n}+n d_{n-1} for n≥2n \geq 2. Deduce the result of part (a).

(c)

Given that exactly mm integers retain their original positions, find the probability that the integer 1 remains in first place.

Problem 3.4.10

In a lecture audience, there are nn students born in 2011, and they were born on independent, uniformly distributed days. Calculate the mean of the number BB of pairs of students sharing a birthday, and show that E[B]>1\mathbb {E}\left[B\right] > 1 if and only if n≥28n \geq 28. Compare this with the result of Problem (1.8.30). Find the variance of BB.

?
Problem 3.4.11

Show that any set of 10 points in the plane R2\mathbb {R}^{2} can be covered by a suitable placement of disjoint open unit disks. [Hint: Consider an infinite array of unit disks whose centres form a triangular lattice.]

?
Problem 3.4.12

Days are either wet or dry, and, given today's weather, tomorrow's is the same as today's with probability pp, and different otherwise. Let wnw_{n} be the probability that the weather nn days into the future from today will be wet. Show that wn+1=1−p+(2p−1)wn−1w_{n+1} = 1-p+(2 p-1) w_{n-1}, and find wnw_{n}. What is the mean number of wet days in the next week?

?
Problem 3.4.13

An urn contains bb balls of which gg are green. Balls are sampled from the urn at random, one by one. After a ball is sampled, its colour is noted, and it is discarded. Find the mean and variance of the number of green balls in a sample of size n(≤b)n( \leq b).

?
Problem 3.4.14

Let n(≥2)n( \geq 2) heterosexual couples be seated randomly at a circular table, subject only to the rule that the sexes alternate. There is no requirement that couples sit together. Let XX be the number of couples seated adjacently. Show that E[X]=2\mathbb {E}\left[X\right] = 2 and Var⁡[X]=2−2/(n−1)\operatorname {Var}\left[X\right] = 2-2 /(n-1).

?
Problem 3.5.1

Each trial may result in any of tt given outcomes, the ii th outcome having probability pip_{i}. Let NiN_{i} be the number of occurrences of the ii th outcome in nn independent trials. Show that

P(Ni=ni for 1≤i≤t)=n!n1!n2!⋯nt!p1n1p2n2⋯ptnt \mathbb {P}\left(N_{i} = n_{i} \text{ for } 1 \leq i \leq t\right) = \frac{n!}{n_{1}!n_{2}!\cdots n_{t}!} p_{1}^{n_{1}} p_{2}^{n_{2}} \cdots p_{t}^{n_{t}}

for any collection n1,n2,…,ntn_{1}, n_{2}, \ldots , n_{t} of non-negative integers with sum nn. The vector NN is said to have the multinomial distribution.

?
Problem 3.5.2

In your pocket is a random number NN of coins, where NN has the Poisson distribution with parameter λ\lambda. You toss each coin once, with heads showing with probability pp each time. Show that the total number of heads has the Poisson distribution with parameter λp\lambda p.

?
Problem 3.5.3

Let XX be Poisson distributed where P(X=n)=pn(λ)=λne−λ/n\mathbb {P}\left(X = n\right) = p_{n}(\lambda ) = \lambda^{n} e^{-\lambda } / n ! for n≥0n \geq 0. Show that P(X≤k)=1−∫0λpk(x)dx\mathbb {P}\left(X \leq k\right) = 1-\int_{0}^{\lambda } p_{k}(x) d x.

?
Problem 3.5.4

A population of bb animals has had a number aa of its members captured, marked, and released. Let XX be the number of animals it is necessary to recapture (without re-release) in order to obtain mm marked animals. Show that

P(X=n)=ab(a−1m−1)(b−an−m)/(b−1n−1) \mathbb {P}\left(X = n\right) = \frac{a}{b}\binom {a-1}{m-1}\binom {b-a}{n-m} /\binom {b-1}{n-1}

and find E[X]\mathbb {E}\left[X\right]. This distribution has been called negative hypergeometric.

?
Problem 3.5.5

Let Λ\Lambda be a positive random variable with density function ff and distribution function FF, and let YY have the Poisson distribution with parameter Λ\Lambda. Show for n=0,1,2,…n = 0,1,2, \ldots that

P(Y≤n)=∫0∞pn(λ)F(λ)dλ,P(Y>n)=∫0∞pn(λ)[1−F(λ)]dλ \mathbb {P}\left(Y \leq n\right) = \int _{0}^{\infty } p_{n}(\lambda ) F(\lambda ) d \lambda , \quad \mathbb {P}\left(Y > n\right) = \int _{0}^{\infty } p_{n}(\lambda )[1-F(\lambda )] d \lambda

where pn(λ)=e−λλn/n!p_{n}(\lambda ) = e^{-\lambda } \lambda^{n} / n!.

?
Problem 3.6.1

Show that the collection of random variables on a given probability space and having finite variance forms a vector space over the reals.

?
Problem 3.6.2

Find the marginal mass functions of the multinomial distribution of Exercise (3.5.1).

?
Problem 3.6.3

Let XX and YY be discrete random variables with joint mass function

f(x,y)=C(x+y−1)(x+y)(x+y+1),x,y=1,2,3,… f(x, y) = \frac{C}{(x+y-1)(x+y)(x+y+1)}, \quad x, y = 1,2,3, \ldots

Find the marginal mass functions of XX and YY, calculate CC, and also the covariance of XX and YY.

?
Problem 3.6.4

Let XX and YY be discrete random variables with mean 00, variance 11, and covariance ρ\rho. Show that E[max⁡{X2,Y2}]≤1+1−ρ2\mathbb {E}\left[\max \left\{ X^{2}, Y^{2}\right\} \right] \leq 1+\sqrt{1-\rho^{2}}.

?
Problem 3.6.5

Let XX and YY be discrete random variables with joint mass function ff.

?
(a)

Show that E[log⁡fX(X)]≥E[log⁡fY(X)]\mathbb {E}\left[\log f_{X}(X)\right] \geq \mathbb {E}\left[\log f_{Y}(X)\right].

(b)

Show that the mutual information

I=E[log⁡(f(X,Y)fX(X)fY(Y))] I = \mathbb {E}\left[\log \left(\frac{f(X, Y)}{f_{X}(X) f_{Y}(Y)}\right)\right]

satisfies I≥0I \geq 0, with equality if and only if XX and YY are independent.

Problem 3.6.6

Let X,Y,ZX, Y, Z be discrete random variables with the property that their values are distinct with probability 1. Let a=P(X>Y),b=P(Y>Z),c=P(Z>X)a = \mathbb {P}\left(X > Y\right), b = \mathbb {P}\left(Y > Z\right), c = \mathbb {P}\left(Z > X\right).

?
(a)

Show that min⁡{a,b,c}≤23\min \left\{ a, b, c\right\} \leq \frac{2}{3}, and give an example where this bound is attained.

(b)

Show that, if X,Y,ZX, Y, Z are independent and identically distributed, then a=b=c=12a = b = c = \frac{1}{2}.

(c)

Find min⁡{a,b,c}\min \left\{ a, b, c\right\} and sup⁡pmin⁡{a,b,c}\sup_{p} \min \left\{ a, b, c\right\} when P(X=0)=1\mathbb {P}\left(X = 0\right) = 1, and Y,ZY, Z are independent with P(Z=1)=P(Y=−1)=p,P(Z=−2)=P(Y=2)=1−p\mathbb {P}\left(Z = 1\right) = \mathbb {P}\left(Y = -1\right) = p, \mathbb {P}\left(Z = -2\right) = \mathbb {P}\left(Y = 2\right) = 1-p. Here, sup⁡p\sup_{p} denotes the supremum as pp varies over [0,1][0,1].

Remark.
?

Part (a) is related to de Condorcet's observation that, in an election, it is possible for more than half of the voters to prefer candidate A to candidate B, more than half B to C, and more than half C to A.

Problem 3.6.7

If one picks a numerical entry at random from an almanac, or the annual accounts of a corporation, the first two significant digits, XX, YY, are found to have approximately the joint mass function

f(x,y)=log⁡10(1+110x+y),1≤x≤9,0≤y≤9. f(x, y) = \log _{10}\left(1+\frac{1}{10 x+y}\right), \quad 1 \leq x \leq 9,0 \leq y \leq 9.

Find the mass function of XX and an approximation to its mean.

?
Remark.
?

A heuristic explanation for this phenomenon may be found in the second of Feller's volumes published in 1971. See also Berger and Hill 2015.

Problem 3.6.8

Let XX and YY have joint mass function

f(j,k)=c(j+k)aj+kj!k!,j,k≥0 f(j, k) = \frac{c(j+k) a^{j+k}}{j!k!}, \quad j, k \geq 0

where aa is a constant. Find c,P(X=j),P(X+Y=r)c, \mathbb {P}\left(X = j\right), \mathbb {P}\left(X+Y = r\right), and E[X]\mathbb {E}\left[X\right].

?
Problem 3.6.9

Let X,Y,ZX, Y, Z be non-degenerate and independent random variables. By considering U=X+Y,V=Y+Z,W=Z−XU = X+Y, V = Y+Z, W = Z-X, or otherwise, show that having positive correlation is not a transitive relation.

?
Problem 3.6.10

Use the identity a2d2+b2c2−2abcd=(ad−bc)2a^{2} d^{2}+b^{2} c^{2}-2 a b c d = (a d-b c)^{2} to prove the Cauchy-Schwarz inequality.

?
Problem 3.6.11

Show that

P(X−E[X]>t)≤Var⁡[X]t2+Var⁡[X],t>0 \mathbb {P}\left(X-\mathbb {E}\left[X\right] > t\right) \leq \frac{\operatorname {Var}\left[X\right]}{t^{2}+\operatorname {Var}\left[X\right]}, \quad t > 0
?
Problem 3.7.1

Show the following:

?
(a)

E[aY+bZ∣X]=aE[Y∣X]+bE[Z∣X]\mathbb {E}\left[a Y+b Z \mid X\right] = a \mathbb {E}\left[Y \mid X\right]+b \mathbb {E}\left[Z \mid X\right] for a,b∈Ra, b \in \mathbb {R},

(b)

E[Y∣X]≥0\mathbb {E}\left[Y \mid X\right] \geq 0 if Y≥0Y \geq 0,

(c)

E[1∣X]=1\mathbb {E}\left[1 \mid X\right] = 1,

(d)

if XX and YY are independent then E[Y∣X]=E[Y]\mathbb {E}\left[Y \mid X\right] = \mathbb {E}\left[Y\right],

(e)

('pull-through property') E[Yg(X)∣X]=g(X)E[Y∣X]\mathbb {E}\left[Y g(X) \mid X\right] = g(X) \mathbb {E}\left[Y \mid X\right] for any suitable function gg,

(f)

('tower property') E[E[Y∣X,Z]∣X]=E[Y∣X]=E[E[Y∣X]∣X,Z]\mathbb {E}\left[\mathbb {E}\left[Y \mid X, Z\right] \mid X\right] = \mathbb {E}\left[Y \mid X\right] = \mathbb {E}\left[\mathbb {E}\left[Y \mid X\right] \mid X, Z\right].

Problem 3.7.2

Suppose that XX and YY are discrete random variables, and that ϕ(X)\phi (X) and ψ(X)\psi (X) are two functions of XX satisfying

E[ϕ(X)g(X)]=E[ψ(X)g(X)]=E[Yg(X)] \mathbb {E}\left[\phi (X) g(X)\right] = \mathbb {E}\left[\psi (X) g(X)\right] = \mathbb {E}\left[Y g(X)\right]

for any function gg for which all the expectations exist. Show that ϕ(X)\phi (X) and ψ(X)\psi (X) are almost surely equal, in that P(ϕ(X)=ψ(X))=1\mathbb {P}\left(\phi (X) = \psi (X)\right) = 1.

?
Problem 3.7.3

Suppose that the conditional expectation of YY given XX is defined as the (almost surely) unique function ψ(X)\psi (X) such that E[ψ(X)g(X)]=E[Yg(X)]\mathbb {E}\left[\psi (X) g(X)\right] = \mathbb {E}\left[Y g(X)\right] for all functions gg for which the expectations exist. Show (a)-(f) of Exercise (3.7.1) above (with the occasional addition of the expression 'with probability 11').

?
Problem 3.7.4

How should we define Var⁡[Y∣X]\operatorname {Var}\left[Y \mid X\right], the conditional variance of YY given XX? Show that Var⁡[Y]=E[Var⁡[Y∣X]]+Var⁡[E[Y∣X]]\operatorname {Var}\left[Y\right] = \mathbb {E}\left[\operatorname {Var}\left[Y \mid X\right]\right]+\operatorname {Var}\left[\mathbb {E}\left[Y \mid X\right]\right].

?
Problem 3.7.5

The lifetime of a machine (in days) is a random variable TT with mass function ff. Given that the machine is working after tt days, what is the mean subsequent lifetime of the machine when:

?
(a)

f(x)=(N+1)−1f(x) = (N+1)^{-1} for x∈{0,1,…,N}x \in \left\{ 0,1, \ldots , N\right\},

(b)

f(x)=2−xf(x) = 2^{-x} for x=1,2,…x = 1,2, \ldots (The first part of Problem (3.11.13) may be useful.)

Problem 3.7.6

Let X1,X2,…X_{1}, X_{2}, \ldots be identically distributed random variables with mean μ\mu and variance σ2\sigma^{2}, and let NN be a random variable taking values in the non-negative integers and independent of the XiX_{i}. Let S=X1+X2+⋯+XNS = X_{1}+X_{2}+\cdots +X_{N}. Show that E[S∣N]=μN\mathbb {E}\left[S \mid N\right] = \mu N, and deduce that E[S]=μE[N]\mathbb {E}\left[S\right] = \mu \mathbb {E}\left[N\right]. Find Var⁡[S]\operatorname {Var}\left[S\right] in terms of the first two moments of NN, using the conditional variance formula of Exercise (3.7.4).

?
Problem 3.7.7

A factory has produced nn robots, each of which is faulty with probability ϕ\phi. To each robot a test is applied which detects the fault (if present) with probability δ\delta. Let XX be the number of faulty robots, and YY the number detected as faulty. Assuming the usual independence, show that

E[X∣Y]={nϕ(1−δ)+(1−ϕ)Y}/(1−ϕδ) \mathbb {E}\left[X \mid Y\right] = \left\{ n \phi (1-\delta )+(1-\phi ) Y\right\} /(1-\phi \delta )
?
Problem 3.7.8

Each child is equally likely to be male or female, independently of all other children.

?
(a)

Show that, in a family of predetermined size, the expected number of boys equals the expected number of girls. Was the assumption of independence necessary?

(b)

A randomly selected child is male; does the expected number of his brothers equal the expected number of his sisters? What happens if you do not require independence?

Problem 3.7.9

Let XX and YY be independent with mean μ\mu. Explain the error in the following equation:

′E[X∣X+Y=z]=E[X∣X=z−Y]=E[z−Y]=z−μ′ { }^{\prime } \mathbb {E}\left[X \mid X+Y = z\right] = \mathbb {E}\left[X \mid X = z-Y\right] = \mathbb {E}\left[z-Y\right] = z-\mu ^{\prime }
?
Problem 3.7.10

A coin shows heads with probability pp. Let XnX_{n} be the number of flips required to obtain a run of nn consecutive heads. Show that E[Xn]=∑k=1np−k\mathbb {E}\left[X_{n}\right] = \sum_{k = 1}^{n} p^{-k}.

?
Problem 3.7.11

Give a definition of the conditional covariance Cov⁡[X,Y∣Z]\operatorname {Cov}\left[X, Y \mid Z\right]. Show that

Cov⁡[X,Y]=E[cov⁡(X,Y∣Z)]+Cov⁡[E[X∣Z],E[Y∣Z]] \operatorname {Cov}\left[X, Y\right] = \mathbb {E}\left[\operatorname {cov}(X, Y \mid Z)\right]+\operatorname {Cov}\left[\mathbb {E}\left[X \mid Z\right], \mathbb {E}\left[Y \mid Z\right]\right]
?
Problem 3.7.12

An urn contains initially bb blue balls and rr red balls, where b,r≥2b, r \geq 2. Balls are drawn one by one without replacement. Show that the mean number of draws until the first colour drawn is first repeated equals 3.

?
Problem 3.7.13
?
(a)

Let XX be uniformly distributed on {0,1,…,n}\left\{ 0,1, \ldots , n\right\}. Show that Var⁡[X]=112n(n+2)\operatorname {Var}\left[X\right] = \frac{1}{12} n(n+2).

(b)

A student sits two examinations, gaining XX and YY marks, respectively. In the interests of economy and fairness, the examiner determines that XX shall be uniformly distributed on {0,1,…,n}\left\{ 0,1, \ldots , n\right\}, and that, conditional on X=k,YX = k, Y shall have the binomial bin⁡(n,k/n)\operatorname {bin}(n, k / n) distribution.

(i) Show that E[Y]=12n\mathbb {E}\left[Y\right] = \frac{1}{2} n, and Var⁡[Y]=112(n2+4n−2)\operatorname {Var}\left[Y\right] = \frac{1}{12}\left(n^{2}+4 n-2\right).

(ii) Find E[X+Y]\mathbb {E}\left[X+Y\right] and Var⁡[X+Y]\operatorname {Var}\left[X+Y\right].

Problem 3.7.14

Let X,YX, Y be discrete integer-valued random variables with the joint mass function

f(x,y)=λye−2λx!(y−x)!,0≤x≤y<∞ f(x, y) = \frac{\lambda ^{y} e^{-2 \lambda }}{x!(y-x)!}, \quad 0 \leq x \leq y < \infty

Show that XX and YY are each Poisson-distributed, and that the conditional distribution of XX given YY is binomial.

?
Problem 3.8.1

Let XX and YY be independent variables, XX being equally likely to take any value in {0,1,…,m}\left\{ 0,1, \ldots , m\right\}, and YY similarly in {0,1,…,n}\left\{ 0,1, \ldots , n\right\}. Find the mass function of Z=X+YZ = X+Y. The random variable ZZ is said to have the trapezoidal distribution.

?
Problem 3.8.2

Let XX and YY have the joint mass function

f(x,y)=C(x+y−1)(x+y)(x+y+1),x,y=1,2,3,… f(x, y) = \frac{C}{(x+y-1)(x+y)(x+y+1)}, \quad x, y = 1,2,3, \ldots

Find the mass functions of U=X+YU = X+Y and V=X−YV = X-Y.

?
Problem 3.8.3

Let XX and YY be independent geometric random variables with respective parameters α\alpha and β\beta. Show that

P(X+Y=z)=αβα−β{(1−β)z−1−(1−α)z−1} \mathbb {P}\left(X+Y = z\right) = \frac{\alpha \beta }{\alpha -\beta }\left\{ (1-\beta )^{z-1}-(1-\alpha )^{z-1}\right\}
?
Problem 3.8.4

Let {Xr:1≤r≤n}\left\{ X_{r}: 1 \leq r \leq n\right\} be independent geometric random variables with parameter pp. Show that Z=∑r=1nXrZ = \sum_{r = 1}^{n} X_{r} has a negative binomial distribution.

?
Problem 3.8.5

Sam rolls 6n6 n dice once; he needs at least nn sixes. Isaac rolls 6(n+1)6(n+1) dice; he needs at least n+1n+1 sixes. Who is more likely to obtain the number of sixes he needs?

?
Problem 3.8.6

Let NN be Poisson distributed with parameter λ\lambda. Show that, for any function gg such that the expectations exist, E[Ng(N)]=λE[g(N+1)]\mathbb {E}\left[N g(N)\right] = \lambda \mathbb {E}\left[g(N+1)\right]. More generally, if S=∑r=1NXrS = \sum_{r = 1}^{N} X_{r}, where {Xr:r≥0}\left\{ X_{r}: r \geq 0\right\} are independent identically distributed non-negative integer-valued random variables, show that

E[Sg(S)]=λE[g(S+X0)X0] \mathbb {E}\left[S g(S)\right] = \lambda \mathbb {E}\left[g\left(S+X_{0}\right) X_{0}\right]
?
Problem 3.8.7

Let S=∑i=1NXiS = \sum_{i = 1}^{N} X_{i}, where the Xi,i≥1X_{i}, i \geq 1, are independent, identically distributed random variables with mean μ\mu and variance σ2\sigma^{2}, and NN is positive, integer-valued, and independent of the XiX_{i}. Show that E[S]=μE[N]\mathbb {E}\left[S\right] = \mu \mathbb {E}\left[N\right], and

Var⁡[S]=σ2E[N]+μ2Var⁡[N] \operatorname {Var}\left[S\right] = \sigma ^{2} \mathbb {E}\left[N\right]+\mu ^{2} \operatorname {Var}\left[N\right]
?
Problem 3.8.8

Let XX and YY be independent random variables with the geometric distributions fX(k)=pqk−1f_{X}(k) = p q^{k-1}, fY(k)=λμk−1f_{Y}(k) = \lambda \mu^{k-1}, for k≥1k \geq 1, where p+q=λ+μ=1p+q = \lambda +\mu = 1 and q≠μq \neq \mu. Write down P(X+Y=n+1,X=k)\mathbb {P}\left(X+Y = n+1, X = k\right), and hence find the distribution of X+YX+Y, and the conditional distribution of XX given that X+Y=n+1X+Y = n+1. Does anything special occur when q=μq = \mu?

?
Problem 3.9.1

Let TT be the time which elapses before a simple random walk is absorbed at either of the absorbing barriers at 0 and NN, having started at kk where 0≤k≤N0 \leq k \leq N. Show that P(T<∞)=1\mathbb {P}\left(T < \infty \right) = 1 and E[Tk]<∞\mathbb {E}\left[T^{k}\right] < \infty for all k≥1k \geq 1.

?
Problem 3.9.2

For simple random walk SS with absorbing barriers at 0 and NN, let WW be the event that the particle is absorbed at 0 rather than at NN, and let pk=P(W∣S0=k)p_{k} = \mathbb {P}\left(W \mid S_{0} = k\right). Show that, if the particle starts at kk where 0<k<N0 < k < N, the conditional probability that the first step is rightwards, given WW, equals ppk+1/pkp p_{k+1} / p_{k}. Deduce that the mean duration JkJ_{k} of the walk, conditional on WW, satisfies the equation

ppk+1Jk+1−pkJk+(pk−ppk+1)Jk−1=−pk, for 0<k<N p p_{k+1} J_{k+1}-p_{k} J_{k}+\left(p_{k}-p p_{k+1}\right) J_{k-1} = -p_{k}, \quad \text{ for } 0 < k < N

subject to the convention that pNJN=0p_{N} J_{N} = 0. Show that we may take as boundary condition J0=0J_{0} = 0. Find JkJ_{k} in the symmetric case, when p=12p = \frac{1}{2}.

?
Problem 3.9.3

With the notation of Exercise (3.9.2), suppose further that at any step the particle may remain where it is with probability rr where p+q+r=1p+q+r = 1. Show that JkJ_{k} satisfies

ppk+1Jk+1−(1−r)pkJk+qpk−1Jk−1=−pk p p_{k+1} J_{k+1}-(1-r) p_{k} J_{k}+q p_{k-1} J_{k-1} = -p_{k}

and that, when ρ=q/p≠1\rho = q / p \neq 1,

Jk=1p−q⋅1ρk−ρN{k(ρk+ρN)−2NρN(1−ρk)1−ρN} J_{k} = \frac{1}{p-q} \cdot \frac{1}{\rho ^{k}-\rho ^{N}}\left\{ k\left(\rho ^{k}+\rho ^{N}\right)-\frac{2 N \rho ^{N}\left(1-\rho ^{k}\right)}{1-\rho ^{N}}\right\}
?
Problem 3.9.4

A coin is tossed repeatedly, heads turning up with probability pp on each toss. Player A wins the game if mm heads appear before nn tails have appeared, and player B wins otherwise. Let pmnp_{m n} be the probability that A wins the game. Set up a difference equation for the pmnp_{m n}. What are the boundary conditions?

?
Problem 3.9.5

Consider a simple random walk on the set {0,1,2,…,N}\left\{ 0,1,2, \ldots , N\right\} in which each step is to the right with probability pp or to the left with probability q=1−pq = 1-p. Absorbing barriers are placed at 0 and NN. Show that the number XX of positive steps of the walk before absorption satisfies

E[X]=12{Dk−k+N(1−pk)} \mathbb {E}\left[X\right] = \frac{1}{2}\left\{ D_{k}-k+N\left(1-p_{k}\right)\right\}

where DkD_{k} is the mean number of steps until absorption having started at kk, and pkp_{k} is the probability of absorption at 0.

?
Problem 3.9.6

Let DkD_{k} be the duration of a random walk on {0,1,2,…,a}\left\{ 0,1,2, \ldots , a\right\} with absorbing barriers at 0 and aa, and started at kk, with steps XiX_{i} satisfying

P(Xi=1)=P(Xi=−1)=p,P(Xi=0)=1−2p \mathbb {P}\left(X_{i} = 1\right) = \mathbb {P}\left(X_{i} = -1\right) = p, \quad \mathbb {P}\left(X_{i} = 0\right) = 1-2 p
?
(a)

When p=12p = \frac{1}{2}, show that

Var⁡[Dk]=13k(a−k){(a−k)2+k2−2} \operatorname {Var}\left[D_{k}\right] = \frac{1}{3} k(a-k)\left\{ (a-k)^{2}+k^{2}-2\right\}
(b)

Deduce (without lengthy calculation) that, for p<12p < \frac{1}{2},

Var⁡[Dk]=k(a−k)(2p)2[1−2p+13{(a−k)2+k2−2}] \operatorname {Var}\left[D_{k}\right] = \frac{k(a-k)}{(2 p)^{2}}\left[1-2 p+\frac{1}{3}\left\{ (a-k)^{2}+k^{2}-2\right\} \right]
Problem 3.9.7

Returns and visits by random walk. Consider a simple symmetric random walk on the set {0,1,2,…,a}\left\{ 0,1,2, \ldots , a\right\} with absorbing barriers at 0 and aa, and starting at kk where 0<k<a0 < k < a. Let rkr_{k} be the probability the walk ever returns to kk, and let vxv_{x} be the mean number of visits to point xx before absorption. Find rkr_{k}, and hence show that

vx={2x(a−k)/a for 0<x<k,2k(a−x)/a for k<x<a v_{x} = \begin{cases} 2 x(a-k) / a & \text{ for } 0 < x < k, \\ 2 k(a-x) / a & \text{ for } k < x < a\end{cases}
?
Problem 3.9.8
?
(a)

"Millionaires should always gamble, poor men never" [J. M. Keynes].

(b)

"If I wanted to gamble, I would buy a casino" [P. Getty].

(c)

"That the chance of gain is naturally overvalued, we may learn from the universal success of lotteries" [Adam Smith, 1776].

Discuss.

Problem 3.10.1

Consider a symmetric simple random walk SS with S0=0S_{0} = 0. Let T=min⁡{n≥1:Sn=0}T = \min \left\{ n \geq 1: S_{n} = 0\right\} be the time of the first return of the walk to its starting point. Show that

P(T=2n)=12n−1(2nn)2−2n \mathbb {P}\left(T = 2 n\right) = \frac{1}{2 n-1}\binom {2 n}{n} 2^{-2 n}

and deduce that E[Tα]<∞\mathbb {E}\left[T^{\alpha }\right] < \infty if and only if α<12\alpha < \frac{1}{2}. You may need Stirling's formula: nn ! ∼nn+12e−n2π\sim n^{n+\frac{1}{2}} e^{-n} \sqrt{2 \pi }

?
Problem 3.10.2

For a symmetric simple random walk starting at 00, show that the mass function of the maximum satisfies P(Mn=r)=P(Sn=r)+P(Sn=r+1)\mathbb {P}\left(M_{n} = r\right) = \mathbb {P}\left(S_{n} = r\right)+\mathbb {P}\left(S_{n} = r+1\right) for r≥0r \geq 0.

?
Problem 3.10.3

For a symmetric simple random walk starting at 00, show that the probability that the first visit to S2nS_{2 n} takes place at time 2k2 k equals the product P(S2k=0)P(S2n−2k=0)\mathbb {P}\left(S_{2 k} = 0\right) \mathbb {P}\left(S_{2 n-2 k} = 0\right), for 0≤k≤n0 \leq k \leq n.

?
Problem 3.10.4

A simple random walk on the integers Z\mathbb {Z} moves one step rightwards with probability pp and otherwise one step leftwards, where p∈(0,1)p \in (0,1). Suppose it starts at 0 and has absorbing barriers at ±a\pm a. Show that the time and place of absorption are independent.

?
Problem 3.10.5

Let {Xm:m≥1}\left\{ X_{m}: m \geq 1\right\} be independent, identically distributed random variables taking integer values such that P(X1≥−1)=1\mathbb {P}\left(X_{1} \geq -1\right) = 1. Let SnS_{n} be the (generalized) random walk given by Sn=k+X1+X2+⋯+XnS_{n} = k+X_{1}+X_{2}+\cdots +X_{n}, where k≥0k \geq 0 is given, and let T=inf⁡{n≥0:Sn=0}T = \inf \left\{ n \geq 0: S_{n} = 0\right\} be the hitting time of 0.

Show by induction that P(T=n)=(k/n)P(Sn=0)\mathbb {P}\left(T = n\right) = (k / n) \mathbb {P}\left(S_{n} = 0\right) when n≥1,k≥0n \geq 1, k \geq 0.

?
Problem 3.11.1
?
(a)

Let XX and YY be independent discrete random variables, and let g,h:R→Rg, h: \mathbb {R} \rightarrow \mathbb {R}. Show that g(X)g(X) and h(Y)h(Y) are independent.

(b)

Show that two discrete random variables XX and YY are independent if and only if fX,Y(x,y)=fX(x)fY(y)f_{X, Y}(x, y) = f_{X}(x) f_{Y}(y) for all x,y∈Rx, y \in \mathbb {R}.

(c)

More generally, show that XX and YY are independent if and only if fX,Y(x,y)f_{X, Y}(x, y) can be factorized as the product g(x)h(y)g(x) h(y) of a function of xx alone and a function of yy alone.

Problem 3.11.2

Show that if Var⁡[X]=0\operatorname {Var}\left[X\right] = 0 then XX is almost surely constant; that is, there exists a∈Ra \in \mathbb {R} such that P(X=a)=1\mathbb {P}\left(X = a\right) = 1. (First show that if E[X2]=0\mathbb {E}\left[X^{2}\right] = 0 then P(X=0)=1\mathbb {P}\left(X = 0\right) = 1.)

?
Problem 3.11.3
?
(a)

Let XX be a discrete random variable and let g:R→Rg: \mathbb {R} \rightarrow \mathbb {R}. Show that, when the sum is absolutely convergent,

E[g(X)]=∑xg(x)P(X=x) \mathbb {E}\left[g(X)\right] = \sum _{x} g(x) \mathbb {P}\left(X = x\right)
(b)

If XX and YY are independent and g,h:R→Rg, h: \mathbb {R} \rightarrow \mathbb {R}, show that E[g(X)h(Y)]=E[g(X)]E[h(Y)]\mathbb {E}\left[g(X) h(Y)\right] = \mathbb {E}\left[g(X)\right] \mathbb {E}\left[h(Y)\right] whenever these expectations exist.

Problem 3.11.4

Let Ω={ω1,ω2,ω3}\Omega = \left\{ \omega_{1}, \omega_{2}, \omega_{3}\right\}, with P(ω1)=P(ω2)=P(ω3)=13\mathbb {P}\left(\omega_{1}\right) = \mathbb {P}\left(\omega_{2}\right) = \mathbb {P}\left(\omega_{3}\right) = \frac{1}{3}. Define X,Y,Z:Ω→RX, Y, Z: \Omega \rightarrow \mathbb {R} by

X(ω1)=1,X(ω2)=2,X(ω3)=3,Y(ω1)=2,Y(ω2)=3,Y(ω3)=1,Z(ω1)=2,Z(ω2)=2,Z(ω3)=1. \begin{aligned} & X\left(\omega _{1}\right) = 1, \quad X\left(\omega _{2}\right) = 2, \quad X\left(\omega _{3}\right) = 3, \\ & Y\left(\omega _{1}\right) = 2, \quad Y\left(\omega _{2}\right) = 3, \quad Y\left(\omega _{3}\right) = 1, \\ & Z\left(\omega _{1}\right) = 2, \quad Z\left(\omega _{2}\right) = 2, \quad Z\left(\omega _{3}\right) = 1. \end{aligned}

Show that XX and YY have the same mass functions. Find the mass functions of X+Y,XYX+Y, X Y, and X/YX / Y. Find the conditional mass functions fY∣Zf_{Y \mid Z} and fZ∣Yf_{Z \mid Y}.

?
Problem 3.11.5

For what values of kk and α\alpha is ff a mass function, where:

?
(a)

f(n)=k/{n(n+1)},n=1,2,…f(n) = k /\left\{ n(n+1)\right\} , n = 1,2, \ldots,

(b)

f(n)=knα,n=1,2,…f(n) = k n^{\alpha }, n = 1,2, \ldots (zeta or Zipf distribution)?

Problem 3.11.6

Let XX and YY be independent Poisson variables with respective parameters λ\lambda and μ\mu. Show that:

?
(a)

X+YX+Y is Poisson, parameter λ+μ\lambda +\mu,

(b)

the conditional distribution of XX, given X+Y=nX+Y = n, is binomial, and find its parameters.

Problem 3.11.7

If XX is geometric, show that P(X=n+k∣X>n)=P(X=k)\mathbb {P}\left(X = n+k \mid X > n\right) = \mathbb {P}\left(X = k\right) for k,n≥1k, n \geq 1. Why do you think that this is called the 'lack of memory' property? Does any other distribution on the positive integers have this property?

?
Problem 3.11.8

Show that the sum of two independent binomial variables, bin⁡(m,p)\operatorname {bin}(m, p) and bin⁡(n,p)\operatorname {bin}(n, p) respectively, is bin⁡(m+n,p)\operatorname {bin}(m+n, p).

?
Problem 3.11.9

Let NN be the number of heads occurring in nn tosses of a biased coin. Write down the mass function of NN in terms of the probability pp of heads turning up on each toss. Prove and utilize the identity

∑i(n2i)x2iyn−2i=12{(x+y)n+(y−x)n} \sum _{i}\binom {n}{2 i} x^{2 i} y^{n-2 i} = \frac{1}{2}\left\{ (x+y)^{n}+(y-x)^{n}\right\}

in order to calculate the probability pnp_{n} that NN is even. Compare with Problem (1.8.20).

?
Problem 3.11.10

An urn contains NN balls, bb of which are blue and r(=N−b)r( = N-b) of which are red. A random sample of nn balls is withdrawn without replacement from the urn. Show that the number BB of blue balls in this sample has the mass function

P(B=k)=(bk)(N−bn−k)/(Nn) \mathbb {P}\left(B = k\right) = \binom {b}{k}\binom {N-b}{n-k} /\binom {N}{n}

This is called the hypergeometric distribution with parameters N,bN, b, and nn. Show further that if N,bN, b, and rr approach ∞\infty in such a way that b/N→pb / N \rightarrow p and r/N→1−pr / N \rightarrow 1-p, then

P(B=k)→(nk)pk(1−p)n−k \mathbb {P}\left(B = k\right) \rightarrow \binom {n}{k} p^{k}(1-p)^{n-k}

You have shown that, for small nn and large NN, the distribution of BB barely depends on whether or not the balls are replaced in the urn immediately after their withdrawal.

?
Problem 3.11.11

Let XX and YY be independent bin⁡(n,p)\operatorname {bin}(n, p) variables, and let Z=X+YZ = X+Y. Show that the conditional distribution of XX given Z=NZ = N is the hypergeometric distribution of Problem (3.11.10).

?
Problem 3.11.12

Suppose XX and YY take values in {0,1}\left\{ 0,1\right\}, with joint mass function f(x,y)f(x, y). Write f(0,0)=af(0,0) = a, f(0,1)=b,f(1,0)=c,f(1,1)=df(0,1) = b, f(1,0) = c, f(1,1) = d, and find necessary and sufficient conditions for XX and YY to be:

?
(a)

uncorrelated,

(b)

independent.

Problem 3.11.13
?
(a)

If XX takes non-negative integer values show that

E[X]=∑n=0∞P(X>n) \mathbb {E}\left[X\right] = \sum _{n = 0}^{\infty } \mathbb {P}\left(X > n\right)
(b)

An urn contains bb blue and rr red balls. Balls are removed at random until the first blue ball is drawn. Show that the expected number drawn is (b+r+1)/(b+1)(b+r+1) /(b+1).

(c)

The balls are replaced and then removed at random until all the remaining balls are of the same colour. Find the expected number remaining in the urn.

(d)

Let XX and YY be independent random variables taking values in the non-negative integers, with finite means. Let U=min⁡{X,Y}U = \min \left\{ X, Y\right\} and V=max⁡{X,Y}V = \max \left\{ X, Y\right\}. Show that

E[U]=∑r=1∞P(X≥r)P(Y≥r)E[V]=∑r=1∞[P(X≥r)+P(Y≥r)−P(X≥r)P(Y≥r)]E[UV]=∑r,s=1∞P(X≥r)P(Y≥s) \begin{aligned} \mathbb {E}\left[U\right] & = \sum _{r = 1}^{\infty } \mathbb {P}\left(X \geq r\right) \mathbb {P}\left(Y \geq r\right) \\ \mathbb {E}\left[V\right] & = \sum _{r = 1}^{\infty }[\mathbb {P}\left(X \geq r\right)+\mathbb {P}\left(Y \geq r\right)-\mathbb {P}\left(X \geq r\right) \mathbb {P}\left(Y \geq r\right)] \\ \mathbb {E}\left[U V\right] & = \sum _{r, s = 1}^{\infty } \mathbb {P}\left(X \geq r\right) \mathbb {P}\left(Y \geq s\right) \end{aligned}
(e)

Let XX take values in the non-negative integers. Show that

E[X2]=E[X]+2∑r=0∞rP(X>r)=∑r=0∞(2r+1)P(X>r) \mathbb {E}\left[X^{2}\right] = \mathbb {E}\left[X\right]+2 \sum _{r = 0}^{\infty } r \mathbb {P}\left(X > r\right) = \sum _{r = 0}^{\infty }(2 r+1) \mathbb {P}\left(X > r\right)

and find a similar formula for E[X3]\mathbb {E}\left[X^{3}\right].

Problem 3.11.14

Let X1,X2,…,XnX_{1}, X_{2}, \ldots , X_{n} be independent random variables, and suppose that XkX_{k} is Bernoulli with parameter pkp_{k}. Show that Y=X1+X2+⋯+XnY = X_{1}+X_{2}+\cdots +X_{n} has mean and variance given by

E[Y]=∑1npk,Var⁡[Y]=∑1npk(1−pk) \mathbb {E}\left[Y\right] = \sum _{1}^{n} p_{k}, \quad \operatorname {Var}\left[Y\right] = \sum _{1}^{n} p_{k}\left(1-p_{k}\right)

Show that, for E[Y]\mathbb {E}\left[Y\right] fixed, Var⁡[Y]\operatorname {Var}\left[Y\right] is a maximum when p1=p2=⋯=pnp_{1} = p_{2} = \cdots = p_{n}. That is to say, the variation in the sum is greatest when individuals are most alike. Is this contrary to intuition?

?
Problem 3.11.15

Let X=(X1,X2,…,Xn)\mathbf{X} = \left(X_{1}, X_{2}, \ldots , X_{n}\right) be a vector of random variables. The covariance matrix V(X)\mathbf{V}(\mathbf{X}) of X\mathbf{X} is defined to be the symmetric nn by nn matrix with entries (vij:1≤i,j≤n)\left(v_{i j}: 1 \leq i, j \leq n\right) given by vij=Cov⁡[Xi,Xj]v_{i j} = \operatorname {Cov}\left[X_{i}, X_{j}\right]. Show that ∣V(X)∣=0\left|\mathbf{V}(\mathbf{X})\right| = 0 if and only if the XiX_{i} are linearly dependent with probability one, in that P(a1X1+a2X2+⋯+anXn=b)=1\mathbb {P}\left(a_{1} X_{1}+a_{2} X_{2}+\cdots +a_{n} X_{n} = b\right) = 1 for some a\mathbf{a} and bb. (∣V∣(\left|\mathbf{V}\right| denotes the determinant of V\mathbf{V}.)

?
Problem 3.11.16

Let XX and YY be independent Bernoulli random variables with parameter 12\frac{1}{2}. Show that X+YX+Y and ∣X−Y∣\left|X-Y\right| are dependent though uncorrelated.

?
Problem 3.11.17

A secretary drops nn matching pairs of letters and envelopes down the stairs, and then places the letters into the envelopes in a random order. Use indicators to show that the number XX of correctly matched pairs has mean and variance 1 for all n≥2n \geq 2. Show that the mass function of XX converges to a Poisson mass function as n→∞n \rightarrow \infty.

?
Problem 3.11.18

Let X=(X1,X2,…,Xn)\mathbf{X} = \left(X_{1}, X_{2}, \ldots , X_{n}\right) be a vector of independent random variables each having the Bernoulli distribution with parameter pp. Let f:{0,1}n→Rf:\left\{ 0,1\right\}^{n} \rightarrow \mathbb {R} be increasing, which is to say that f(x)≤f(y)f(\mathbf{x}) \leq f(\mathbf{y}) whenever xi≤yix_{i} \leq y_{i} for each ii.

?
(a)

Let e(p)=E[f(X)]e(p) = \mathbb {E}\left[f(\mathbf{X})\right]. Show that e(p1)≤e(p2)e\left(p_{1}\right) \leq e\left(p_{2}\right) if p1≤p2p_{1} \leq p_{2}.

(b)

FKG inequality. Let ff and gg be increasing functions from {0,1}n\left\{ 0,1\right\}^{n} into R\mathbb {R}. Show by induction on nn that Cov⁡[f(X),g(X)]≥0\operatorname {Cov}\left[f(\mathbf{X}), g(\mathbf{X})\right] \geq 0.

Problem 3.11.19

Let R(p)R(p) be the reliability function of a network GG with a given source and sink, each edge of which is working with probability pp, and let AA be the event that there exists a working connection from source to sink. Show that

R(p)=∑ωIA(ω)pN(ω)(1−p)m−N(ω) R(p) = \sum _{\omega } I_{A}(\omega ) p^{N(\omega )}(1-p)^{m-N(\omega )}

where ω\omega is a typical realization (i.e. outcome) of the network, N(ω)N(\omega ) is the number of working edges of ω\omega, and mm is the total number of edges of GG.

Deduce that R′(p)=Cov⁡[IA,N]/{p(1−p)}R^{\prime }(p) = \operatorname {Cov}\left[I_{A}, N\right] /\left\{ p(1-p)\right\}, and hence that

R(p)(1−R(p))p(1−p)≤R′(p)≤mR(p)(1−R(p))p(1−p) \frac{R(p)(1-R(p))}{p(1-p)} \leq R^{\prime }(p) \leq \sqrt{\frac{m R(p)(1-R(p))}{p(1-p)}}
?
Problem 3.11.20

Let R(p)R(p) be the reliability function of a network GG, each edge of which is working with probability pp.

?
(a)

Show that R(p1p2)≤R(p1)R(p2)R\left(p_{1} p_{2}\right) \leq R\left(p_{1}\right) R\left(p_{2}\right) if 0≤p1,p2≤10 \leq p_{1}, p_{2} \leq 1.

(b)

Show that R(pγ)≤R(p)γR\left(p^{\gamma }\right) \leq R(p)^{\gamma } for all 0≤p≤10 \leq p \leq 1 and γ≥1\gamma \geq 1.

Problem 3.11.21

In a certain style of detective fiction, the sleuth is required to declare "the criminal has the unusual characteristics...; find this person and you have your man". Assume that any given individual has these unusual characteristics with probability 10−710^{-7} independently of all other individuals, and that the city in question contains 10710^{7} inhabitants. Calculate the expected number of such people in the city.

?
(a)

Given that the police inspector finds such a person, what is the probability that there is at least one other?

(b)

If the inspector finds two such people, what is the probability that there is at least one more?

(c)

How many such people need be found before the inspector can be reasonably confident that he has found them all?

(d)

For the given population, how improbable should the characteristics of the criminal be, in order that he (or she) be specified uniquely?

Problem 3.11.22

In 1710, J. Arbuthnot observed that male births had exceeded female births in London for 82 successive years. Arguing that this showed the two sexes cannot be equally likely, since 2−822^{-82} is very small, he attributed this run of masculinity to Divine Providence. Let us assume that each birth results in a girl with probability p=0.485p = 0.485, and that the outcomes of different confinements are independent of each other. Ignoring the possibility of twins (and so on), show that the probability that girls outnumber boys in 2n2 n live births is no greater than (2nn)pnqn{q/(q−p)}\binom {2 n}{n} p^{n} q^{n}\left\{ q /(q-p)\right\}, where q=1−pq = 1-p. Suppose that 20,000 children are born in each of 82 successive years. Show that the probability that boys outnumber girls every year is at least 0.99. You may need Stirling's formula.

?
Problem 3.11.23

Consider a symmetric random walk with an absorbing barrier at NN and a reflecting barrier at 0 (so that, when the particle is at 00, it moves to 1 at the next step). Let αk(j)\alpha_{k}(j) be the probability that the particle, having started at kk, visits 0 exactly jj times before being absorbed at NN. We make the convention that, if k=0k = 0, then the starting point counts as one visit. Show that

αk(j)=N−kN2(1−1N)j−1,j≥1,0≤k≤N \alpha _{k}(j) = \frac{N-k}{N^{2}}\left(1-\frac{1}{N}\right)^{j-1}, \quad j \geq 1,0 \leq k \leq N
?
Problem 3.11.24

Problem of the points (3.9.4). A coin is tossed repeatedly, heads turning up with probability pp on each toss. Player A wins the game if heads appears at least mm times before tails has appeared nn times; otherwise player B wins the game. Find the probability that A wins the game.

?
Problem 3.11.25

A coin is tossed repeatedly, heads appearing on each toss with probability pp. A gambler starts with initial fortune kk (where 0<k<N0 < k < N); he wins one point for each head and loses one point for each tail. If his fortune is ever 0 he is bankrupted, whilst if it ever reaches NN he stops gambling to buy a Jaguar. Suppose that p<12p < \frac{1}{2}. Show that the gambler can increase his chance of winning by doubling the stakes. You may assume that kk and NN are even.

What is the corresponding strategy if p≥12p \geq \frac{1}{2}?

?
Problem 3.11.26

A compulsive gambler is never satisfied. At each stage he wins £1£ 1 with probability pp and loses £1£ 1 otherwise. Find the probability that he is ultimately bankrupted, having started with an initial fortune of £k£ k.

?
Problem 3.11.27

Let {Xn:n≥1}\left\{ X_{n}: n \geq 1\right\} be independent, identically distributed random variables taking integer values. Let S0=0,Sn=∑i=1nXiS_{0} = 0, S_{n} = \sum_{i = 1}^{n} X_{i}. The range RnR_{n} of S0,S1,…,SnS_{0}, S_{1}, \ldots , S_{n} is the number of distinct values taken by the sequence. Show that P(Rn=Rn−1+1)=P(S1S2⋯Sn≠0)\mathbb {P}\left(R_{n} = R_{n-1}+1\right) = \mathbb {P}\left(S_{1} S_{2} \cdots S_{n} \neq 0\right), and deduce that, as n→∞n \rightarrow \infty,

1nE[Rn]→P(Sk≠0 for all k≥1) \frac{1}{n} \mathbb {E}\left[R_{n}\right] \rightarrow \mathbb {P}\left(S_{k} \neq 0 \text{ for all } k \geq 1\right)

Hence show that, for the simple random walk, n−1E[Rn]→∣p−q∣n^{-1} \mathbb {E}\left[R_{n}\right] \rightarrow \left|p-q\right| as n→∞n \rightarrow \infty.

?
Problem 3.11.28

Arc sine law for maxima. Consider a symmetric random walk SS starting from the origin, and let Mn=max⁡{Si:0≤i≤n}M_{n} = \max \left\{ S_{i}: 0 \leq i \leq n\right\}. Show that, for i=2k,2k+1i = 2 k, 2 k+1, the probability that the walk reaches M2nM_{2 n} for the first time at time ii equals 12P(S2k=0)P(S2n−2k=0)\frac{1}{2} \mathbb {P}\left(S_{2 k} = 0\right) \mathbb {P}\left(S_{2 n-2 k} = 0\right).

?
Problem 3.11.29

Let SS be a symmetric random walk with S0=0S_{0} = 0, and let NnN_{n} be the number of points that have been visited by SS exactly once up to time nn. Show that E[Nn]=2\mathbb {E}\left[N_{n}\right] = 2.

?
Problem 3.11.30

Consider the following fragment of verse entitled 'Note for the scientist'.

People who have three daughters try for more, And then its fifty-fifty they'll have four, Those with a son or sons will let things be, Hence all these surplus women, QED.

?
(a)

What do you think of the argument?

(b)

Show that the mean number of children of either sex in a family whose fertile parents have followed this policy equals 1. (You should assume that each delivery yields exactly one child whose sex is equally likely to be male or female.) Discuss.

Problem 3.11.31

Let β>1\beta > 1, let p1,p2,…p_{1}, p_{2}, \ldots denote the prime numbers, and let N1,N2,…N_1, N_2, \ldots be independent random variables, NiN_i having mass function P(Ni=k)=(1−γi)γik\mathbb {P}\left(N_i = k\right) = \left(1-\gamma_{i}\right) \gamma_{i}^{k} for k≥0k \geq 0, where γi=pi−β\gamma_{i} = p_{i}^{-\beta } for all ii. Show that M=∏i=1∞piNiM = \prod_{i = 1}^{\infty } p_{i}^{N_i} is a random integer with mass function P(M=m)=Cm−β\mathbb {P}\left(M = m\right) = C m^{-\beta } for m≥1m \geq 1 (this may be called the Dirichlet distribution), where CC is a constant satisfying

C=∏i=1∞(1−1piβ)=(∑m=1∞1mβ)−1 C = \prod _{i = 1}^{\infty }\left(1-\frac{1}{p_{i}^{\beta }}\right) = \left(\sum _{m = 1}^{\infty } \frac{1}{m^{\beta }}\right)^{-1}
?
Problem 3.11.32

N+1N+1 plates are laid out around a circular dining table, and a hot cake is passed between them in the manner of a symmetric random walk: each time it arrives on a plate, it is tossed to one of the two neighbouring plates, each possibility having probability 12\frac{1}{2}. The game stops at the moment when the cake has visited every plate at least once. Show that, with the exception of the plate where the cake began, each plate has probability 1/N1 / N of being the last plate visited by the cake.

?
Problem 3.11.33

There are (nm)\binom {n}{m} points ranked in order of merit with no matches. You seek to reach the best, BB. If you are at the jj th best, you step to any one of the j−1j-1 better points, with equal probability of stepping to each. Let rjr_{j} be the expected number of steps to reach BB from the jj th best vertex. Show that rj=∑k=1j−1k−1r_{j} = \sum_{k = 1}^{j-1} k^{-1}. Give an asymptotic expression for the expected time to reach BB from the worst vertex, for large m,nm, n.

?
Problem 3.11.34

There are nn unstable molecules in a row, m1,m2,…,mnm_{1}, m_{2}, \ldots , m_{n}. One of the n−1n-1 pairs of neighbours, chosen at random, combines to form a stable dimer; this process continues until there remain UnU_{n} isolated molecules no two of which are adjacent. Show that the probability that m1m_{1} remains isolated is ∑r=0n−1(−1)r/r!→e−1\sum_{r = 0}^{n-1}(-1)^{r} / r! \rightarrow e^{-1} as n→∞n \rightarrow \infty. Deduce that lim⁡n→∞n−1E[Un]=e−2\lim_{n \rightarrow \infty } n^{-1} \mathbb {E}\left[U_{n}\right] = e^{-2}.

?
Problem 3.11.35

Let {Ir:1≤r≤n}\left\{ I_{r}: 1 \leq r \leq n\right\} be independent Bernoulli random variables with respective parameters {pr:1≤r≤n}\left\{ p_{r}: 1 \leq r \leq n\right\} satisfying pr≤c<1p_{r} \leq c < 1 for all rr and some cc. Let λ=∑r=1npr\lambda = \sum_{r = 1}^{n} p_{r} and X=∑r=1nXrX = \sum_{r = 1}^{n} X_{r}. Show that

P(X=k)=λke−λk!{1+O(λmax⁡rpr+k2λmax⁡pr)} \mathbb {P}\left(X = k\right) = \frac{\lambda ^{k} e^{-\lambda }}{k!}\left\{ 1+\mathrm{O}\left(\lambda \max _{r} p_{r}+\frac{k^{2}}{\lambda } \max p_{r}\right)\right\}
?
Problem 3.11.36

The length of the tail of the rr th member of a troop of NN chimeras is xrx_{r}. A random sample of nn chimeras is taken (without replacement) and their tails measured. Let IrI_{r} be the indicator of the event that the rr th chimera is in the sample. Set

Xr=xrIr,Yˉ=1n∑r=1NXr,μ=1N∑r=1Nxr,σ2=1N∑r=1N(xr−xˉ)2 X_{r} = x_{r} I_{r}, \quad \bar{Y} = \frac{1}{n} \sum _{r = 1}^{N} X_{r}, \quad \mu = \frac{1}{N} \sum _{r = 1}^{N} x_{r}, \quad \sigma ^{2} = \frac{1}{N} \sum _{r = 1}^{N}\left(x_{r}-\bar{x}\right)^{2}

Show that E[Yˉ]=μ\mathbb {E}\left[\bar{Y}\right] = \mu, and Var⁡[Yˉ]=(N−n)σ2/{n(N−1)}\operatorname {Var}\left[\bar{Y}\right] = (N-n) \sigma^{2} /\left\{ n(N-1)\right\}.

?
Problem 3.11.37

Any individual in a group GG contracts a certain disease CC with probability γ\gamma; such individuals are hospitalized with probability cc. Independently of this, anyone in GG may be in hospital with probability aa, for some other reason. Let XX be the number in hospital, and YY the number in hospital who have CC (including those with CC admitted for any other reason). Show that the correlation between XX and YY is

ρ(X,Y)=γp1−γp⋅(1−a)(1−γc)a+γc−aγc \rho (X, Y) = \sqrt{\frac{\gamma p}{1-\gamma p} \cdot \frac{(1-a)(1-\gamma c)}{a+\gamma c-a \gamma c}}

where p=a+c−acp = a+c-a c. It has been stated erroneously that, when ρ(X,Y)\rho (X, Y) is near unity, this is evidence for a causal relation between being in GG and contracting CC.

?
Problem 3.11.38

A telephone sales company attempts repeatedly to sell new kitchens to each of the NN families in a village. Family ii agrees to buy a new kitchen after it has been solicited KiK_{i} times, where the KiK_{i} are independent identically distributed random variables with mass function f(n)=P(Ki=n)f(n) = \mathbb {P}\left(K_{i} = n\right). The value ∞\infty is allowed, so that f(∞)≥0f(\infty ) \geq 0. Let XnX_{n} be the number of kitchens sold at the nnth round of solicitations, so that Xn=∑i=1NI{Ki=n}X_{n} = \sum_{i = 1}^{N} I_{\left\{ K_{i} = n\right\} }. Suppose that NN is a random variable with the Poisson distribution with parameter vv.

?
(a)

Show that the XnX_{n} are independent random variables, XrX_{r} having the Poisson distribution with parameter vf(r)v f(r).

(b)

The company loses heart after the TT th round of calls, where T=inf⁡{n:Xn=0}T = \inf \left\{ n: X_{n} = 0\right\}. Let S=X1+X2+⋯+XTS = X_{1}+X_{2}+\cdots +X_{T} be the number of solicitations made up to time TT. Show further that E[S]=νE[F(T)]\mathbb {E}\left[S\right] = \nu \mathbb {E}\left[F(T)\right] where F(k)=f(1)+f(2)+⋯+f(k)F(k) = f(1)+f(2)+\cdots +f(k).

Problem 3.11.39

A particle performs a random walk on the non-negative integers as follows. When at the point nn (>0)( > 0) its next position is uniformly distributed on the set {0,1,2,…,n+1}\left\{ 0,1,2, \ldots , n+1\right\}. When it hits 0 for the first time, it is absorbed. Suppose it starts at the point aa.

?
(a)

Find the probability that its position never exceeds aa, and prove that, with probability 11, it is absorbed ultimately.

(b)

Find the probability that the final step of the walk is from 1 to 0 when a=1a = 1.

(c)

Find the expected number of steps taken before absorption when a=1a = 1.

Problem 3.11.40

Let GG be a finite graph with neither loops nor multiple edges, and write dvd_{v} for the degree of the vertex vv. An independent set is a set of vertices no pair of which is joined by an edge. Let α(G)\alpha (G) be the size of the largest independent set of GG. Use the probabilistic method to show that α(G)≥∑v1/(dv+1)\alpha (G) \geq \sum_{v} 1 /\left(d_{v}+1\right).

?
Remark.
?

This conclusion is sometimes referred to as Turán's theorem.

Problem 3.11.41

Kelly betting, or proportional investment. A gambler (or 'investor') makes a sequence of bets of the following type: on each bet, for a given stake SS, the return is either the loss of the stake with probability q(=1−p)q( = 1-p) or a win totalling (1+r)S(1+r) S with probability pp. (Assume the usual independence.) The entry fee is cSc S where c<rc < r. Show that the mean gain per play for stake SS is gSg S, where g=pr−q−cg = p r-q-c (with a negative value indicating a loss).

The gambler decides to bet a fixed fraction ff of her current fortune at each stage. That is, given a current fortune FF, she bets fFf F for some fixed ff. Show that her resulting fortune is

F′=F{1+f[(1+r)I−(1+c)]} F^{\prime } = F\left\{ 1+f[(1+r) I-(1+c)]\right\}

where II is the indicator function of a win. Suppose p>(1+c)/(1+r)p > (1+c) /(1+r). The gambler considers two policies for choosing ff : for given FF,

?
(a)

maximize E[F′]\mathbb {E}\left[F^{\prime }\right],

(b)

maximize E[log⁡F′]\mathbb {E}\left[\log F^{\prime }\right]

Find an expression for ff in each case. Show that fa>fbf_{a} > f_{b}, and explain why a cautious gambler may prefer (b) to (a) even though this entails a slower rate of growth of her expected fortune.

Problem 3.11.42

Let x1,x2,…,xrx_{1}, x_{2}, \ldots , x_{r} be given reals with r≥2r \geq 2, and let the sequence {Xn:n≥1}\left\{ X_{n}: n \geq 1\right\} of random variables be given as follows. First, Xn=xnX_{n} = x_{n} for n≤rn \leq r. For n≥rn \geq r, we set Xn+1=XUn+XVnX_{n+1} = X_{U_{n}}+X_{V_{n}} where Un,VnU_{n}, V_{n} are uniformly distributed on {1,2,…,n}\left\{ 1,2, \ldots , n\right\}, and the family {Un,Vn:n≥r}\left\{ U_{n}, V_{n}: n \geq r\right\} are independent. Show that

E[Xn]=2nr(r+1)∑k=1rxk \mathbb {E}\left[X_{n}\right] = \frac{2 n}{r(r+1)} \sum _{k = 1}^{r} x_{k}

Enthusiasts may care to show that, when r=1=x1r = 1 = x_{1} and n→∞n \rightarrow \infty,

1n2E[Xn2]→12πsinh⁡π \frac{1}{n^{2}} \mathbb {E}\left[X_{n}^{2}\right] \rightarrow \frac{1}{2 \pi } \sinh \pi
?
Problem 3.11.43

Let X1=1X_{1} = 1. For n≥1n \geq 1, let Xn+1=XUn−XVnX_{n+1} = X_{U_{n}}-X_{V_{n}} where Un,VnU_{n}, V_{n} are uniformly distributed on {1,2,…,n}\left\{ 1,2, \ldots , n\right\}, and the family {Un,Vn:n≥1}\left\{ U_{n}, V_{n}: n \geq 1\right\} are independent. Show that

1nVar⁡[Xn]→−sin⁡(π3)π3 as n→∞ \frac{1}{n} \operatorname {Var}\left[X_{n}\right] \rightarrow -\frac{\sin (\pi \sqrt{3})}{\pi \sqrt{3}} \quad \text{ as } n \rightarrow \infty

You may find it useful to recall Euler's sine formula:

sin⁡(πx)=πx∏n=1∞(1−x2n2) \sin (\pi x) = \pi x \prod _{n = 1}^{\infty }\left(1-\frac{x^{2}}{n^{2}}\right)
?
Problem 3.11.44

Gollum has concealed the ring of power in a box chosen randomly from a row of n≥1n \geq 1 such boxes. Bilbo opens a box at random. If the ring is not there, his occult powers are sufficient for him to learn whether the ring lies to the right or the left, and he opens further boxes accordingly. Find an expression for the mean number bnb_{n} of boxes opened before finding the ring, and deduce that bn∼2log⁡nb_{n} \sim 2 \log n as n→∞n \rightarrow \infty.

?
Problem 3.11.45

In a variant of the previous problem, we have n=2r−1n = 2^{r}-1, and Bilbo invariably chooses the middle box. Find the mean number mrm_{r} of boxes inspected, and the asymptotics of mrm_{r} as r→∞r \rightarrow \infty.

?
Problem 3.11.46

A bad fairy has cursed you. A good fairy has concealed the magic word that cancels the curse in one of nn numbered boxes, and she has told you that it is in box ii with probability pip_{i}, for i=1,2,…,ni = 1,2, \ldots , n. Each day you are permitted to look in one box.

?
(a)

Assume that, each day, you inspect a box chosen at random, box ii being chosen with probability cic_{i}, and, furthermore, boxes chosen on different days are independent. Find the mean number of days that elapse before your release from the curse, and find the mass function cc that minimizes this mean value.

(b)

Suppose now that you remember the results of your previous failed searches. What now is your optimal policy, and what is the mean number of days that elapse?

(c)

After each search, the bad fairy removes the magic word, which is immediately replaced by the good fairy in an independently chosen box, with the same distribution at each replacement. What now is your optimal policy? Find the mean number of elapsed days.

Problem 3.11.47

Gwen and John play 'best of 2n+12 n+1 games', and play stops as soon as either has won n+1n+1 games. Gwen wins each game with probability γ∈(0,1)\gamma \in (0,1) and John otherwise (with probability δ=1−γ\delta = 1-\gamma). Different games have independent winners. Write down an expression for the probability frf_{r} that Gwen wins rr games altogether, given that John has won the match, and deduce that rfr=(n+r)γfr−1r f_{r} = (n+r) \gamma f_{r-1} for 0<r≤n0 < r \leq n.

Hence or otherwise, prove that the mean total number TnT_{n} of games in the match is

Tn=(n+1)(γ(1−Pn)δ+δPnγ+1)−(2n+1)(2nn)(γδ)n T_{n} = (n+1)\left(\frac{\gamma \left(1-P_{n}\right)}{\delta }+\frac{\delta P_{n}}{\gamma }+1\right)-(2 n+1)\binom {2 n}{n}(\gamma \delta )^{n}

where PnP_{n} is the probability that Gwen wins the match. When p=12p = \frac{1}{2}, show that

Tn=2n−2n/π+2+O(n−1/2), as n→∞ T_{n} = 2 n-2 \sqrt{n / \pi }+2+\mathrm{O}\left(n^{-1 / 2}\right), \quad \text{ as } n \rightarrow \infty
?
Problem 3.11.48

Let S(n,k)S(n, k) be the number of ways to partition N={1,2,…,n}N = \left\{ 1,2, \ldots , n\right\} into kk non-empty parts. Suppose each element of NN is coloured with one of cc distinct colours. Prove that

cn=∑k=1nS(n,k)c(c−1)⋯(c−k+1) c^{n} = \sum _{k = 1}^{n} S(n, k) c(c-1) \cdots (c-k+1)

Deduce that the nnth moment of the Poisson distribution with parameter 1 equals the number bnb_{n} of ways to partition NN, that is, bn=∑k=1nS(n,k)b_{n} = \sum_{k = 1}^{n} S(n, k).

?
Problem 3.11.49

Let β,γ∈[0,1]\beta , \gamma \in [0,1]. Bertha and Harold play a game of rackets. Bertha wins the point with probability β\beta when she serves, and Harold wins with probability γ\gamma when he serves. The first player to win nn points wins the game, and Bertha serves first. Consider the following rules for changing server.

?
(a)

Service alternates between players.

(b)

Service is retained until a point is lost, and then passes to the other player.

(c)

Service is retained until a point is won, and then passes to the other player.

(d)

Bertha serves the first nn points, and then Harold serves any further points required to decide the outcome.

Show that P(Bertha wins)\mathbb {P}\left(\text{Bertha wins}\right) is the same for all four rules.

Problem 3.11.50

Let X,YX, Y be discrete random variables with respective mass functions fX,fYf_{X}, f_{Y}, and joint mass function fX,Yf_{X, Y}. Define the:

 entropy of X:H(X)=−E[log⁡fX(X)] joint entropy of X given Y:H(X,Y)=−E[log⁡fX,Y(X,Y)] conditional entropy of X given Y:H(X∣Y)=−E[log⁡fX∣Y(X∣Y)] \begin{aligned} \text{ entropy of } X & : H(X) = -\mathbb {E}\left[\log f_{X}(X)\right] \\ \text{ joint entropy of } X \text{ given } Y & : H(X, Y) = -\mathbb {E}\left[\log f_{X, Y}(X, Y)\right] \\ \text{ conditional entropy of } X \text{ given } Y & : H(X \mid Y) = -\mathbb {E}\left[\log f_{X \mid Y}(X \mid Y)\right] \end{aligned}
?
(a)

Show that H(X+a)=H(X)H(X+a) = H(X) for a∈Ra \in \mathbb {R}.

(b)

Show that H(X)−H(X∣Y)=I(X;Y)H(X)-H(X \mid Y) = I(X ; Y), where II is the mutual information of Exercise (3.6.5).

(c)

Show when XX and YY are independent that H(X,Y)=H(X)+H(Y)H(X, Y) = H(X)+H(Y).

(d)

Show that the entropy of the binomial distribution bin⁡(n,p)\operatorname {bin}(n, p) is non-decreasing in nn.

(e)

Find the entropy of the geometric distribution, parameter pp, and show it is decreasing in pp.

Remark.
?

It is normal to use logarithms to base 2 in information theory, but we use natural logarithms here.

Problem 3.11.51
?
(a)

Show that the entropy H(λ)H(\lambda ), as defined in Problem (3.11.50), of the Poisson distribution with parameter λ\lambda (using natural logarithms) is given by

H(λ)=λ−λlog⁡λ+e−λ∑m=0∞λmlog⁡(m!)m! H(\lambda ) = \lambda -\lambda \log \lambda +e^{-\lambda } \sum _{m = 0}^{\infty } \frac{\lambda ^{m} \log (m!)}{m!}
(b)

Recalling Exercise (3.6.5) and Example (3.7.5), show that the mutual information of the number NN of hens and the number KK of chicks is I(N;K)=H(λ)−H(λ(1−p))I(N ; K) = H(\lambda )-H(\lambda (1-p)). Deduce that H(λ)H(\lambda ) is increasing in λ\lambda.

(c)

Spend a short time seeking to show the last statement directly from the expression in part (a).

Problem 3.11.52

Let β>1\beta > 1, and let M,LM, L be independent random variables with the Dirichlet distribution with parameter β\beta of Problem (3.11.31).

?
(a)

Show that the events Ep={X is divisible by p}E_{p} = \left\{ X \text{ is divisible by }p\right\} are independent for pp prime.

(b)

Deduce Euler's formula

∏p prime (1−1pβ)=1ζ(β) \prod _{p \text{ prime }}\left(1-\frac{1}{p^{\beta }}\right) = \frac{1}{\zeta (\beta )}

where ζ(β)\zeta (\beta ) is the Riemann zeta function, ζ(β)=∑m=1∞m−β\zeta (\beta ) = \sum_{m = 1}^{\infty } m^{-\beta }, and β>1\beta > 1.

(c)

Show that the probability that MM is 'square-free' (that is, indivisible by any perfect square other than 1) equals 1/ζ(2β)1 / \zeta (2 \beta )

(d)

Let HH be the highest common factor of MM and LL. Prove that

P(H=m)=m−2βζ(2β),m=1,2,… \mathbb {P}\left(H = m\right) = \frac{m^{-2 \beta }}{\zeta (2 \beta )}, \quad m = 1,2, \ldots
Problem 3.11.53

A strict Oxford secret is a secret which you may tell to no more than one other person. Of a group of n+1n+1 Oxonians, one learns a strict secret. In accordance with the rules, she tells it to one other person selected uniformly at random from the rest of the group. Each confidante tells the secret to exactly one person picked at random from the rest of the group, excluding the person from whom they heard the secret. When someone who already knows the secret hears it repeated, the entire process ceases.

Let SS be the total number of people who eventually know the secret. Find the distribution of SS, and show that

1nE[S]→π/2,1nVar⁡[S]→12(4−π), as n→∞ \frac{1}{\sqrt{n}} \mathbb {E}\left[S\right] \rightarrow \sqrt{\pi / 2}, \quad \frac{1}{n} \operatorname {Var}\left[S\right] \rightarrow \frac{1}{2}(4-\pi ), \quad \text{ as } n \rightarrow \infty

With nr‾n^{\underline{r}} denoting n!/(n−r)n!/(n-r) !, you may find it useful that

∑r=1∞nr‾nr∼πn/2,∑r=1∞rnr‾nr∼n \sum _{r = 1}^{\infty } \frac{n^{\underline{r}}}{n^{r}} \sim \sqrt{\pi n / 2}, \quad \sum _{r = 1}^{\infty } r \frac{n^{\underline{r}}}{n^{r}} \sim n
?
Problem 3.11.54

From a pack of nn cards, two different cards are selected randomly, and they are transposed. Let prp_{r} be the probability that any given card (say, the top card) is in its original position after r>0r > 0 such independent transpositions.

?
(a)

Show that

pr=1n+n−1n(n−3n−1)r p_{r} = \frac{1}{n}+\frac{n-1}{n}\left(\frac{n-3}{n-1}\right)^{r}
(b)

Find E[Cr]\mathbb {E}\left[C_{r}\right], where CrC_{r} is the number of cards in their original place after rr random transpositions.

(c)

Show that, for large nn, the number rr of transpositions needed for E[Cr]≈2\mathbb {E}\left[C_{r}\right] \approx 2 is approximately 12nlog⁡n\frac{1}{2} n \log n.

Problem 3.11.55

The dd-cube CdC_{d} is the graph with vertex-set {0,1}d\left\{ 0,1\right\}^{d}, and an edge between two vertices x=(x1,x2,…,xd)\mathbf{x} = \left(x_{1}, x_{2}, \ldots , x_{d}\right) and y=(y1,y2,…,yd)\mathbf{y} = \left(y_{1}, y_{2}, \ldots , y_{d}\right) if and only if ∑i∣xi−yi∣=1\sum_{i}\left|x_{i}-y_{i}\right| = 1. A particle pursues a random walk on CdC_{d}. At each epoch of time, it moves from its current position to a neighbour chosen uniformly at random, with the usual independence. Two vertices are called 'antipodal' if the graph-distance between them equals dd.

Show that the mean first passage time mdm_{d} of the walker between two antipodal vertices satisfies μd∼2d\mu_{d} \sim 2^{d} as d→∞d \rightarrow \infty.

?
Problem 3.11.56
?
(a)

A particle performs a type of random walk on the set S={1,2,…,n}S=\left\{ 1,2, \ldots , n\right\}. Let XrX_{r} be the particle's position at time rr. Given that Xr=x,Xr+1X_{r}=x, X_{r+1} is chosen uniformly at random from the set {1}∪{x+1,x+2,…,n}\left\{ 1\right\} \cup \left\{ x+1, x+2, \ldots , n\right\}. The particle stops moving at the first instant it arrives at either 1 or nn (so that 1 and nn are 'absorbing', but absorption does not occur at time 0 even if X0∈{1,n}X_{0} \in \left\{ 1, n\right\} ). The point m∈{1,2,…,n}m \in \left\{ 1,2, \ldots , n\right\} is said to be hit if Xr=mX_{r}=m for some r≥1r \geq 1. Show that

P(m is hit ∣X0=1)={12 if m=1,1n−m+2 if m≥2 \mathbb {P}\left(m \text{ is hit } \mid X_{0}=1\right) = \begin{cases} \frac{1}{2} & \text{ if } m=1, \\ \frac{1}{n-m+2} & \text{ if } m \geq 2\end{cases}
(b)

The nn passengers on a flight in an airplane with nn seats have been told their seat numbers. They board the plane one by one. The first person sits in a seat chosen uniformly at random from the nn seats then available. Subsequent passengers sit in their assigned seats whenever they find them available, or otherwise in a randomly chosen empty seat. For m≥2m \geq 2, what is the probability that the mm th passenger finds his or her assigned seat to be already occupied?

Problem 3.11.57
?
(a)

Let XX be a random variable with E[X]>0\mathbb {E}\left[X\right]>0 and 0<E[X2]<∞0<\mathbb {E}\left[X^{2}\right]<\infty. Show that

P(X>aE[X])≥(1−a)2E[X]2E[X2] \mathbb {P}\left(X > a \mathbb {E}\left[X\right]\right) \geq \frac{(1-a)^2\mathbb {E}\left[X\right]^2}{\mathbb {E}\left[X^2\right]}
(b)

Deduce that, if P(X≥0)=1\mathbb {P}\left(X \geq 0\right)=1,

P(X=0)≤Var⁡[X]E[X2]≤Var⁡[X]E[X]2 \mathbb {P}\left( X = 0\right) \leq \frac{\operatorname {Var}\left[X\right]}{\mathbb {E}\left[X^2\right]} \leq \frac{\operatorname {Var}\left[X\right]}{\mathbb {E}\left[X\right]^2}
(c)

Let A1,A2,…,AnA_{1}, A_{2}, \ldots , A_{n} be events, and let X=∑r=1n  1ArX=\sum_{r=1}^{n} \; \mathbb {1}_{A_r} be the sum of their indicator functions. Show that

P(X=0)≤1E[X]+1E[X]2∑∗P(Ar∩As) \mathbb {P}\left(X = 0\right) \leq \frac{1}{\mathbb {E}\left[X\right]} + \frac{1}{\mathbb {E}\left[X\right]^2} \sum ^{*} \mathbb {P}\left(A_r \cap A_s\right)

where the summation ∑∗\sum^{*} is over all distinct unordered pairs r,sr, s such that ArA_{r} and AsA_{s} are not independent.